OpenStudy (anonymous):

Write the slope-intercept form of the line −4x + 6y = 3

OpenStudy (anonymous):

y = mx + b 6y = 4x + 3 6y/6 = 4x/6 + 3/6 y = 4x/6 + 1/2

OpenStudy (anonymous):

y= 2/3x+1/2!